if sin A = 5/13 for angle A in quadrant 1, find sin A/2
Answer by stanbon(75887)

if sin A = 5/13 for angle A in quadrant 1, find sin A/2
--------
sin(A/2) = sqrt[(1-cos(A))/2]
---
cos(A) = sqrt[13^2-5^2]/13 = 12/13
-----
Ans: sin(A/2) = sqrt[(1-(12/13))/2] = sqrt[(1/13)/2] = sqrt(1/26)
--------------
Cheers,
Stan H.
------------